OpenStudy (anonymous):

If y varies directly as x, and y is 6 when x is 72, what is the value of y when x is 8?

OpenStudy (anonymous):

1st find the constant

OpenStudy (anonymous):

\[y \alpha x\]

OpenStudy (anonymous):

y=kx

OpenStudy (anonymous):

6=k(72), k=6/72, k=1/12

OpenStudy (anonymous):

y=(1/12)(8) ,y=2/3

OpenStudy (anonymous):

now it makes scense thanks

OpenStudy (anonymous):

welcome^^